Graph the equation
6y-12=0

Answer by Edwin McCravy(20055)

6y-12=0

When you have an equation of a line with only one letter in it
like this, one way to graph it is to add 0 times the missing 
letter to it.  In this case x is missing, so add 0x to the 
equation.  That is, make it:

0x+6y-12=0

Now make a table of values, substituting different arbitrary values
for the missing letter:

 x | y
 -----
 0 |
 1 |
-3 |
 5 |

Substitute x=0:   Substitute x=1:   Substitute x=-3   Substitute x=5 

  0x+6y-12=0        0x+6y-12=0        0x+6y-12=0        0x+6y-12=0
0(0)+6y-12=0      0(1)+6y-12=0     0(-3)+6y-12=0      0(5)+6y-12=0
   0+6y-12=0         0+6y-12=0         0+6y-12=0         0+6y-12=0
     6y-12=0           6y-12=0           6y-12=0           6y-12=0
        6y=12             6y=12             6y=12            6y=12
         y=2               y=2               y=2              y=2

Now we have filled in the table:

 x | y
 -----
 0 | 2
 1 | 2
-3 | 2
 5 | 2

So we plot the points (0,2), (1,2), (-3,2), and (5,2)



Now draw a line through them:



Can you think of an easier way to do this when there is a missing letter
in your equation?
If not, it's OK, for this way will always work.

Edwin